OpenStudy (anonymous):

1. Marco and his two younger sisters would like to purchase a silver charm bracelet for their mother’s birthday. They went to the mall and found what they were looking for at Store A. In Store A, the bracelet without charms costs $85 and each charm costs $15

OpenStudy (imstuck):

What's the question?

OpenStudy (anonymous):

A. Use function notation that models the total price of the bracelet and how that price is based on the number of charms. Explain the reasoning behind your equation. B. What would be a reasonable domain for this function based on this scenario? Explain why this is a proper domain. C. If Marco and his sisters have saved $250, make a graph to show all of the charms they could purchase with a bracelet if they had saved between $85 and $250. D. Marco found five charms that he likes. Using your function, determine the cost of the bracelet he would make. E. As they walked down the hall of the mall, Marco saw this sales flyer for Store B on a bench.

OpenStudy (anonymous):

aight, so where are we at?

OpenStudy (imstuck):

For the first part, we will solve for x (the number of charms) in terms of the cost of the bracelet. So the equation would be c = 85 + 15x, c being the cost, 85 being the base price of the plain bracelet, and 15x being 15 times the number of charms, which is the changing coefficient. For the second part, a reasonable domain depends upon if there is at least one charm. If there is, the least the bracelet could cost would be 85 + 15 up to however many charms are purchased. The domain is "x is greater than or equal to 100". For part c, this is an equation of sorts. If they have 250 to spend on the bracelet, we set c to = 250 and solve for x, which iis the number of charms. 250 = 85 + 15x. Doing the math, x comes out to be 11 charms. If x is 0 when y is 85, and x is 11 when is 250, that is a line on a graph with the x axis being the number of charms and the y axis being the cost at x. See that part? You could draw the graph yourself and see how you could find the cost of the bracelet for any number of charms purchased. For part d: Fill in 5 for x in our equation c = 85 + 15x. Can you solve that? What is the question for the part about the flyer from store B?

OpenStudy (anonymous):

@IMStuck Marco is wondering from which store he should buy the items so that he spends the least amount of money. Which store should Marco use if he wanted the five charms and silver bracelet all together? What advice can you tell Marco to help him shop wisely depending on the number of charms he wants to buy? Justify how you reasoned your answers

OpenStudy (anonymous):

sorry

OpenStudy (imstuck):

That's ok. So now we have to do some comparative shopping! The equation for store A is c = 85 + 15x. The equation for store B is c = 100 + 12x. x represents the number of charms in both cases. Fill in 5 for x in the first equation to get c = 85 + 15*5. That gives you 160. In the second equation, fill in 5 for x to get c = 100 + 12*5. That gives you 160 also. So it seems that no matter where Marco goes, he will spend $160. Tell him that, and also he should go where the charms are nicer or where the salespeople are nicer, because cost is equal. See all that?

OpenStudy (anonymous):

thanks, and this isn't a real scenario, it was in my assessment @IMStuck

OpenStudy (anonymous):

he was telling you what to tell him bc it says in your question to give him adivce lmao @KryoWolf